OpenStudy (anonymous):

URGENTTT A student scored 84 and 87 on her first two quizzes. Write and solve a compound inequality to find the possible values for a third quiz score that would give her an average between 85 and 90, inclusive. -0.5

OpenStudy (anonymous):

B

OpenStudy (anonymous):

the answer is 87

OpenStudy (nottim):

Should give an explanation; I actually wanna know how to do this.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

The student already scored 84 and 87, so the average between the scores would be \[\frac{ 84 + 87 + n }{ 3 }\] and n is the third quiz score, so, the average must be between -0.5 and 4.5 inclusive The "inclusive" means less or equal than So \[85<\frac{ 84 + 87 + n }{ 3 }\le90\] If you multiply the entire inequality by 3 you get \[255<84+87+n \le 270\] and now you can sum up the 84+87 which is 171 now, you substract 171 to the entire inequality to leave x alove so \[255-171<171+n-171 \le 270-171\] \[84<n \le 99\]

OpenStudy (anonymous):

Correction: "the average must be between 80 and 90 inclusive"